Please answer the following 3 questions showing work and explaining
media%2F1ea%2F1ea24b33-8340-47b1-a64f-95
Answer #1

It is a right-tailed test.

The correct choice of the hypothesis is

10.c)H0 : p1 = p2 , HA: P1> P2

11)pooled proportion= 7 ni + n2 = (84+28)/(120+80) = 0.560

Ans:-a)0.560

12) z = P1-P2 = 0.7-0.35 0.45(1-0.45) + = 0.35/0.0718=4.874

Ans:-d)4.874.